# CPCL: Cross-Modal Prototypical Contrastive Learning for Weakly Supervised Text-based Person Re-Identification

### v1

@@ 3.4 Outlier Pseudo Label Mining
3.4.1 The Refined Learning Stage
@@
For example, in $OPLM_{v\rightarrow t}$, for an outlier image sample $v^{o}_{i}$, we can find the corresponding text description according to the known image text corresponding index. However, it should be noted that there may be more than one paired text for an image. Therefore, its paired captions are
@@
where $P^{t}=\{t_{1},\ldots,t_{p}\}$ represents that there are $p$ pieces of caption paired with this image, and $PIS(\cdot)$ denotes the process of paired instance searching in another modality. Subsequently, we proceed to filter $P^{t}$ by eliminating the outlier samples and obtain $P^{t}_{\textrm{filter}}$. If $||P^{t}_{\textrm{filter}}||=0$, $v^{o}_{i}$ remains as a outlier sample. Conversely, if $P^{t}_{\textrm{filter}}=\{t_{1},\ldots,t_{k}\}$, it indicates that there are $k$ clustered samples. We then process all the samples in the collection $P^{t}_{\textrm{filter}}$ to identify their corresponding images:
$P^{v}=\{PIS(t_{i})\mid t_{i}\in P^{t}_{\textrm{filter}}\ ,i=1,\ldots,k\}\ ,$ (10)
where, $P^{v}=\{v_{1},\ldots,v_{k}\}$ denotes $k$ images corresponding to $P^{t}_{\textrm{filter}}$ in a one-to-one manner. Following this, $P^{v}$ undergoes a filtering process where outlier samples are removed, resulting in $P^{v}_{\textrm{filter}}$. If $||P^{v}_{\textrm{filter}}||=0$, then $v^{o}_{i}$ is kept as a outlier sample. However, if $P^{v}_{\textrm{filter}}=\{v_{1},\ldots,v_{q}\}$, it signifies the presence of $q$ clustered image samples. Subsequently, the distance from $v^{o}_{i}$ to all samples in $P^{v}_{\textrm{filter}}$ is calculated, and the nearest sample is identified as $v_{i}$
$v_{i}=\mathop{\arg\max}\limits_{v_{i}\in P^{v}_{\textrm{filter}},v_{i}\not\in U^{v}}\langle v^{o}_{i}\cdot v_{i}\rangle,i=1,\ldots,q\ ,$ (11)
@@
If $v_{i}$ is still an outlier sample, it will be added to the collection $U^{v}$. The process will then return to formula Eq. 11 for re-selection until $P^{v}_{\textrm{filter}}=U^{v}$. If the outlier sample $v^{o}_{i}$ remains unassigned to a specific category, it will be discarded, and the algorithm will proceed to the next outlier sample. During the process of Outlier Pseudo Label Mining (OPLM), all the outlier samplers in both modalities are traversed.
3.4.2 The Supplementary Learning Stage
At the supplementary learning stage, the image-text pairs that are not mined by the refined learning stage compose the training dataset. The main objective loss function is ITC loss. For each image feature $f^{v}_{i}$, we obtain an InfoNCEOord et al. (2018) loss between its image feature $f^{v}_{i}$ and all text features $f^{t}_{j}$ in the batch,
$\mathcal{L}_{\textrm{i2t}}^{itc}=-\log\frac{\exp\left(f^{v}_{i}\cdot f^{t}_{i}/\tau\right)}{\sum_{k=1}^{N}\exp\left(f^{v}_{i}\cdot f^{t}_{k}/\tau\right)},$ (13)
where the $N$ denotes the batch size, and $\tau$ denotes the hyperparameter of temperature. Symmetrically, for each text feature, the InfoNCE loss is denoted as $\mathcal{L}_{\textrm{t2i}}^{itc}$. The total loss function for ITC loss is defined as:
$\mathcal{L}_{\textrm{itc}}=\mathcal{L}_{\textrm{i2t}}^{itc}+\mathcal{L}_{\textrm{t2i}}^{itc}.$ (14)
@@ 3.4.2 The Supplementary Learning Stage
Method Type Ref Image Enc. Text Enc. Rank-1 Rank-5 Rank-10 mAP mINP
Fully Supervised Text-based Person Re-Identification
DSSLZhu et al. (2021) L MM21 RN50 BERT 59.98 80.41 87.56 - -
SSANDing et al. (2021) L arXiv21 RN50 LSTM 61.37 80.15 86.73 - -
LBULWang et al. (2022b) L MM22 RN50 BERT 64.04 82.66 87.22 - -
Han et al.Han et al. (2021) G BMVC21 CLIP-RN101 CLIP-Xformer 64.08 81.73 88.19 60.08 -
CAIBCWang et al. (2022a) L MM22 RN50 BERT 64.43 82.87 88.37 - -
AXM-NetFarooq et al. (2022) L MM22 RN50 BERT 64.44 80.52 86.77 58.73 -
IVTShu et al. (2022) G ECCVW22 ViT-Base BERT 65.59 83.11 89.21 - -
CFineYan et al. (2022) L arXiv22 CLIP-ViT BERT 69.57 85.93 91.15 - -
IRRAJiang and Ye (2023) G CVPR23 CLIP-ViT CLIP-Xformer 73.38 89.93 93.71 66.13 50.24
RaSaBai et al. (2023) G IJCAI23 ALBEF-ViT ALBEF-BERT 76.51 90.29 94.25 69.38 -
Weakly Supervised Text-based Person Re-Identification
CMMTZhao et al. (2021) G ICCV21 ResNet LSTM 57.10 78.14 85.23 - -
CMMT(CLIP-ViT-B/16)* G ICCV21 CLIP-ViT CLIP-Xformer 59.57 79.53 86.53 54.66 39.78
CAIBCWang et al. (2022a) G MM22 ResNet BERT 58.64 79.02 85.93 - -
Baseline (CLIP-ViT-B/16) G - CLIP-ViT CLIP-Xformer 58.45 78.87 85.30 54.14 39.83
CPCL (Ours) G - CLIP-ViT CLIP-Xformer 70.03 87.28 91.78 63.19 47.54
Table 1: Performance comparisons with SOTA methods on CUHK-PEDES. “G” and “L” correspond to the global-matching and local-matching methods, respectively. (*) the implementation is based on the CMMT’s paper.
4 Experiment
We conduct experiments on three TPRe-ID datasets: CUHK-PEDESLi et al. (2017b), ICFG-PEDESDing et al. (2021) and RSTPReidZhu et al. (2021). The introduction of each dataset and evaluation metrics are shown in the Appendix.
Method Type Rank-1 Rank-5 Rank-10 mAP mINP
Fully Supervised Text-based Person Re-Identification
Dual PathZheng et al. (2020) G 38.99 59.44 68.41 - -
ViTAAWang et al. (2020) L 50.98 68.79 75.78 - -
SSANDing et al. (2021) L 54.23 72.63 79.53 - -
IVTShu et al. (2022) G 56.04 73.60 80.22 - -
CFineYan et al. (2022) L 60.83 76.55 82.42 - -
IRRAJiang and Ye (2023) G 63.46 80.25 85.82 38.06 7.93
RaSaBai et al. (2023) G 65.28 80.40 85.12 41.29 -
Weakly Supervised Text-based Person Re-Identification
CMMT(CLIP-ViT-B/16)* G 54.27 71.17 77.86 33.17 5.73
Baseline (CLIP-ViT-B/16) G 53.83 73.33 80.46 30.60 5.35
CPCL (Ours) G 62.60 79.07 84.46 36.16 6.31
@@ 4 Experiment
Table 2: Performance comparisons with SOTA methods on ICFG-PEDES dataset.
Method Type Rank-1 Rank-5 Rank-10 mAP mINP
Fully Supervised Text-based Person Re-Identification
DSSLZhu et al. (2021) G 39.05 62.60 73.95 - -
SSANDing et al. (2021) L 43.50 67.80 77.15 - -
LBULWang et al. (2022b) L 45.55 68.20 77.85 - -
IVTShu et al. (2022) G 46.70 70.00 78.80 - -
CFineYan et al. (2022) L 50.55 72.50 81.60 - -
IRRAJiang and Ye (2023) G 60.20 81.30 88.20 47.17 25.28
RaSaBai et al. (2023) G 66.90 86.50 91.35 52.31 -
Weakly Supervised Text-based Person Re-Identification
CMMT(CLIP-ViT-B/16)* G 52.25 76.45 84.55 41.98 22.00
Baseline (CLIP-ViT-B/16) G 53.10 75.70 83.60 38.65 21.06
CPCL (Ours) G 58.35 81.05 87.65 45.81 23.87
@@
Table 3: Performance comparisons with SOTA methods on RSTPReid dataset.
No. Methods Components CUHK-PEDES
OPLM $\mathcal{L}_{pcm}$ $\mathcal{L}_{icpm}$ Rank-1 Rank-5 Rank-10
0 Baseline 58.45 78.87 85.30
1 +OPLM+$\mathcal{L}_{id}$ 61.44 81.48 88.56
2 +$\mathcal{L}_{pcm}$ 65.14 84.62 90.51
3 +$\mathcal{L}_{icpm}$ 65.08 83.38 89.30
4 +$\mathcal{L}_{pcm}$+$\mathcal{L}_{icpm}$ 68.76 86.76 91.23
5 +OPLM+$\mathcal{L}_{icpm}$ 67.99 86.08 91.63
6 +OPLM+$\mathcal{L}_{pcm}$ 68.24 86.06 91.26
7 CPCL 70.03 87.28 91.78
@@
No. Methods Components CUHK-PEDES
$PCM_{c}$ Memory update Two-stage OPLM Rank-1 Rank-5 Rank-10
0 $PCM_{s}$ 68.24 86.06 91.23
1 no update 69.10 86.18 91.19
2 One-stage OPLM 68.76 86.11 91.26
3 CPCL 70.03 87.28 91.78
@@
4.1 Implementation Details
We adopt a pretrained CLIP-ViT-B/16Radford et al. (2021) as the backbone. The batch size is set to 128, and the model is trained with Adam optimizerKingma and Ba (2015) for 60 epochs with a learning rate initialized to $1\times 10^{-5}$ and cosine learning rate decay. In the beginning, the 5 warm-up epochs linearly increase the learning rate from $1\times 10^{-6}$ to $1\times 10^{-5}$. Following the clustering-based method CMMTZhao et al. (2021), we use DBSCANEster et al. (1996) for clustering before each epoch. Note that we utilize the CLIP-ViT-B/16 model fine-tuned with ITC loss as the Baseline. More details can be found in Appendix.
4.2 Comparison with State-of-the-art Methods
As CMMTZhao et al. (2021) is only tested on the CUHK-PEDES dataset and does not public its training code, to conduct the comparison of the backbone and additional module between CMMT and our method, we implemented CMMT with CLIP as the backbone to access its performance on ICFG-PEDES and RSTPeid.
4.2.1 Comparison with Weakly supervised TPRe-ID Methods
We compared our method with prior methods on three datasets, as shown in Table 1, Table 2 and Table 3, respectively. Our method CPCL achieves the SOTA performance in terms of all metrics, outperforming existing methods by a large margin. Specifically, compared with the Baseline, CPCL gains a significant Rank@1 improvement of 11.58%, 8.77% and 5.25% on the three datasets, respectively.
4.2.2 Comparison with Fully Supervised TPRe-ID Methods
To see the gaps between weakly and fully supervised TPRe-ID, we also list the performance of the fully supervised TPRe-ID methods in the 1st group on three datasets, as shown in Table 1, Table 2 and Table 3 respectively. Compared with current RaSaJiang and Ye (2023), CPCL is lower than 3.35%, 0.86% and 1.85% in Rank@1 accuracy on CUHK-PEDES, ICFG-PEDES and RSTPReid datasets, respectively. While there is a gap between our method and the current fully supervised SOTA method(RaSaBai et al. (2023)), our method is already very close to some fully supervised methods.
@@ 4.3 Ablation Study
In this chapter, we thoroughly analyze the effectiveness of each module in the CPCL method as shown in Table 4.
@@
4.3.1 Prototypical Multi-modal Memory (PMM) and Hybrid-level Cross-modal Matching (HCM)
@@
It should be noted that PMM and HCM need to be combined to work properly. HCM comprises two components: PCM and ICPM. As shown in Table 4, No.0 vs No.2, No.0 vs No.3, No.1 vs No.6, and No.1 vs No.5, we readily observe that the inclusion of the PMM and HCM proves beneficial for enhancing the model’s performance. Removing this module results in a certain degree of performance degradation.
4.3.2 Outlier Pseudo Label Mining (OPLM)
@@
As shown in Table 4, No.0 vs No.1, No.2 vs No.6 and No.3 vs No.5, we conducted an ablation study of OPLM. OPLM makes significant contributions to the model’s performance. Being a plug-and-play component, OPLM can seamlessly integrate into weakly supervised TPRe-ID methods which are based on clustering algorithms. The removal of this module results in the inability to extract valuable diverse samples, leading to performance degradation of the model. More analysis of OPLM is shown in the Appendix.
4.3.3 cross-modadl PCM vs single-modal PCM
We attempt two types of Prototypical Contrastive Matching (PCM) to enforce the embedding of a sample to be more similar to its corresponding prototypes compared to other prototypes: cross-modal PCM and single-modal PCM, denoted as $PCM_{c}$ and $PCM_{s}$. The experimental results are shown No.0 vs No.3 in Table 5. When employing $PCM_{s}$, it achieves competitive performance on CUHK-PEDES. $PCM_{c}$ surpasses $PCM_{s}$ in terms of all metrics and we argue that $PCM_{c}$ promotes alignment between visual and textual samples. In conclusion, experiments prove that $PCM_{c}$ is more suitable for model learning.
4.3.4 Exploration of Momentum updating scheme for PMM
We employ a momentum updating scheme for PMM to update the prototypes and conduct experiments to highlight the necessity of a momentum updating scheme for PMM and its impact on the performance of CPCL, as illustrated in No.1 vs No.3 in Table 5. Disabling the momentum updating scheme eliminates prototypical feature variance and instability in the training process caused by update issues. However, it also hinders the model’s further performance. We argue that the momentum update of prototypical features enhances our method’s performance and adaptability. This is achieved by enabling the model to effectively capture changes and evolution in the data while preserving overall stability.

### v2

@@ 3.4 Outlier Pseudo Label Mining
The Refined Learning Stage
@@
For example, in $OPLM_{v\rightarrow t}$, for an outlier image sample $v^{o}_{i}$, we can find the corresponding text description according to image text corresponding index. However, it should be noted that there may be more than one paired text for one image. Therefore, its paired captions are
@@
where $P^{t}=\{t_{1},\ldots,t_{p}\}$ represents that there are $p$ captions paired with this image, and $PIS(\cdot)$ denotes the process of paired instance searching in text modality. Subsequently, we filter $P^{t}$ by eliminating outlier text samples and obtain $P^{t}_{filter}$. If $||P^{t}_{filter}||=0$, $v^{o}_{i}$ remains as an outlier image sample. In contrast, if $P^{t}_{filter}=\{t_{1},\ldots,t_{k}\}$, it indicates that there are $k$ clustered samples. We then process all samples in the collection $P^{t}_{filter}$ to identify their corresponding images:
$P^{v}=\{PIS(t_{i})\mid t_{i}\in P^{t}_{filter}\ ,i=1,\ldots,k\}\ ,$ (16)
where, $P^{v}=\{v_{1},\ldots,v_{k}\}$ denotes $k$ images corresponding to $P^{t}_{filter}$ in a one-to-one manner. Following this, $P^{v}$ undergoes a filtering process in which outlier image samples are removed, resulting in $P^{v}_{filter}$. If $||P^{v}_{filter}||=0$, then $v^{o}_{i}$ is kept as an outlier image sample. Subsequently, the distance from $v^{o}_{i}$ to all samples in $P^{v}_{filter}$ is calculated, and the nearest sample is identified as $v_{i}$
$v_{i}=\mathop{\arg\max}\limits_{v_{i}\in P^{v}_{filter},v_{i}\not\in U^{v}}\langle v^{o}_{i}\cdot v_{i}\rangle,i=1,\ldots,q\ ,$ (17)
@@
Method R@1 R@5 R@10 mAP mINP
Fully Supervised Text-based Person Retrieval
PDG(Wu et al. 2024) 69.47 87.13 92.13 60.56 -
IRRA(Jiang and Ye 2023) 73.38 89.93 93.71 66.13 50.24
RaSa(Bai et al. 2023) 76.51 90.29 94.25 69.38 -
APTM(Yang et al. 2023) 76.53 90.04 94.15 66.91 -
CFAM(Zuo et al. 2024) 76.71 91.83 95.96 68.47 -
Tan et al.(Tan et al. 2024) 76.82 - - 69.55 -
DualFocus(Deng et al. 2024) 77.43 90.73 94.20 68.35 53.56
MARS(Ergasti et al. 2024) 77.62 90.63 94.27 71.41 -
BAMG(Cheng et al. 2024) 79.98 92.31 94.03 68.55 -
Weakly Supervised Text-based Person Retrieval
CMMT(Zhao et al. 2021) 57.10 78.14 85.23 - -
CMMT(CLIP-ViT-B/16)* 59.57 79.53 86.53 54.66 39.78
CAIBC(Wang et al. 2022) 58.64 79.02 85.93 - -
ECCA(Gong et al. 2024b) 68.13 87.26 91.88 - -
CCL(Du et al. 2024) 67.25 86.10 91.45 60.83 -
CANC(Gong et al. 2024a) 69.61 88.10 91.47 - -
CPCL(CLIP-RN50) 60.09 80.72 87.70 53.47 36.57
CPCL(CLIP-RN101) 64.88 84.02 89.70 57.76 41.22
CPCL(CLIP-ViT-B/16) 70.03 87.28 91.78 63.19 47.54
Table 1: Performance comparisons with SOTA methods on CUHK-PEDES. “G” and “L” correspond to the global-matching and local-matching methods, respectively. (*) denotes our implementation based on the CMMT’s paper.
The Supplementary Learning Stage
At the supplementary learning stage, the image-text pairs that are not mined by the refined learning stage compose the training dataset. The main objective loss function is ITC loss. For each image feature $f^{v}_{i}$, we obtain an InfoNCE(Oord et al. 2018) loss between its image feature $f^{v}_{i}$ and all text features $f^{t}_{j}$ in the batch,
$\mathcal{L}_{v2t}^{itc}=-\log\frac{\exp\left(f^{v}_{i}\cdot f^{t}_{i}/\tau\right)}{\sum_{k=1}^{N}\exp\left(f^{v}_{i}\cdot f^{t}_{k}/\tau\right)},$ (19)
where the $N$ denotes the batch size, and $\tau$ denotes the hyperparameter of temperature. Symmetrically, for each text feature, the InfoNCE loss is denoted as $\mathcal{L}_{t2i}^{itc}$. The total loss function for ITC loss is defined as:
$\mathcal{L}_{itc}=\mathcal{L}_{v2t}^{itc}+\mathcal{L}_{t2v}^{itc}.$ (20)
@@
4 Experiments
4.1 Evaluation Metrics and Dataset
For all our experiments, we utilize recall at Rank-k (k=1, 5, 10) metric to evaluate the performance of the retrieval system. Additionally, for a comprehensive evaluation, we incorporate the mean Average Precision (mAP) and mean Inverse Negative Penalty (mINP)(Ye et al. 2021) as additional retrieval criteria. For dataset, we use CUHK-PEDES, ICDG-PEDES, and RSTPReid. For more information on them, please refer to Appendix G.
Method R@1 R@5 R@10 mAP mINP
Fully Supervised Text-based Person Retrieval
PDG(Wu et al. 2024) 57.69 75.79 82.67 36.07 -
IRRA(Jiang and Ye 2023) 63.46 80.25 85.82 38.06 7.93
RaSa(Bai et al. 2023) 65.28 80.40 85.12 41.29 -
APTM(Yang et al. 2023) 68.51 82.09 87.56 41.22 -
CFAM(Zuo et al. 2024) 66.58 82.47 87.37 40.46 -
DualFocus(Deng et al. 2024) 67.87 81.93 87.13 40.13 9.14
Tanet al.(Tan et al. 2024) 67.05 - - 41.51 -
MARS(Ergasti et al. 2024) 67.60 81.47 85.79 44.93 -
BAMG(Cheng et al. 2024) 71.70 86.34 89.71 42.37 -
Weakly Supervised Text-based Person Retrieval
ECCA(Gong et al. 2024b) 58.89 77.46 83.34 - -
CCL(Du et al. 2024) 58.33 76.75 83.38 32.66 -
CANC(Gong et al. 2024a) 60.52 78.36 84.13 - -
CPCL(CLIP-RN50) 52.20 72.45 79.77 28.82 4.12
CPCL(CLIP-RN101) 55.15 74.38 81.24 30.47 4.47
CPCL(CLIP-ViT-B/16) 62.60 79.07 84.46 36.16 6.31
@@ 4.1 Evaluation Metrics and Dataset
Table 2: Performance comparisons with methods on ICFG-PEDES dataset. (*) denotes our implementation based on the CMMT’s paper
Method R@1 R@5 R@10 mAP mINP
Fully Supervised Text-based Person Retrieval
IRRA(Jiang and Ye 2023) 60.20 81.30 88.20 47.17 25.28
RaSa(Bai et al. 2023) 66.90 86.50 91.35 52.31 -
APTM(Yang et al. 2023) 67.50 85.70 91.45 52.56 -
CFAM(Zuo et al. 2024) 63.54 84.75 92.32 50.48 -
DualFocus(Deng et al. 2024) 69.12 86.68 92.31 52.55 27.87
Tanet al.(Tan et al. 2024) 68.50 - - 53.02 -
MARS(Ergasti et al. 2024) 67.55 86.65 91.35 52.92 -
BAMG(Cheng et al. 2024) 69.73 87.65 93.33 55.21 -
Weakly Supervised Text-based Person Retrieval
CMMT(CLIP-ViT-B/16)* 52.25 76.45 84.55 41.98 22.00
ECCA(Gong et al. 2024b) 58.49 80.75 85.60 - -
CCL(Du et al. 2024) 51.30 75.25 84.60 41.10 -
CANC(Gong et al. 2024a) 56.24 80.15 86.61 - -
CPCL(CLIP-RN50) 49.55 75.85 84.15 37.37 16.22
CPCL(CLIP-RN101) 52.60 78.30 85.45 39.14 17.06
CPCL(CLIP-ViT-B/16) 58.35 81.05 87.65 45.81 23.87
@@
Table 3: Performance comparisons with methods on RSTPReid dataset. (*) denotes our implementation based on the CMMT’s paper
4.2 Implementation Details
Motivated by the successful use of CLIP knowledge to text-image person retrieval, we directly initialize our CPCL with the full CLIP model to enhance our method crucial cross-modal alignment capabilities. More details about the implementation, please infer to Appendix F.
4.3 Comparison with State-of-the-art Methods
As CMMT(Zhao et al. 2021) is only tested on the CUHK-PEDES dataset, we implemented CMMT with CLIP as the backbone to access its performance on ICFG-PEDES and RSTPeid. We also conduct domain generalization experiments to assess the generalization ability of CPCL which are reported in the Appendix D due to space limitations.
Comparison with Weakly supervised Text-based Person Retrieval Methods
We compared our method with prior methods on three datasets, as shown in Table 1, Table 2 and Table 3, respectively. Our method CPCL achieves the SOTA performance in terms of all metrics, outperforming existing methods by a large margin. Specifically, compared with the SOTA, CPCL gains a significant Rank-1 improvement of 0.42%, 2.08% and 2.11% on the three datasets, respectively.
Comparison with Fully Supervised Text-based Person Retrieval Methods
To see the gaps between weakly and fully supervised text-based person retrieval, we also list the performance of the fully supervised text-based person retrieval methods in the 1st group on three datasets, as shown in Table 1, Table 2 and Table 3 respectively. Our method remains highly competitive even in comparison to fully supervised text-based person retrieval approaches, surpassing several of them in performance.
No. Components CUHK-PEDES
OPLM $\mathcal{L}_{pcm}$ $\mathcal{L}_{icpm}$ R@1 R@5 R@10
0 - - - 58.45 78.87 85.30
1 - - 61.44 81.48 88.56
2 - - 65.14 84.62 90.51
3 - - 65.08 83.38 89.30
4 - 68.76 86.76 91.23
5 - 67.99 86.08 91.63
6 - 68.24 86.06 91.26
7 70.03 87.28 91.78
@@ 4.3 Comparison with State-of-the-art Methods
No. Components CUHK-PEDES
$PCM_{c}$
Memory
update
Two-stage
OPLM
R@1 R@5 R@10
0 - 68.24 86.06 91.23
1 - 69.10 86.18 91.19
2 - 68.76 86.11 91.26
3 70.03 87.28 91.78
@@ 4.4 Ablation Study
In this section, we thoroughly analyze the effectiveness of each module in the CPCL. The overall results are reported in Table 4.
@@
Prototypical Multi-modal Memory (PMM) and Hybrid-level Cross-modal Matching (HCM)
@@
In contrast to existing CLIP-based methods that solely establish one-to-one connections between image-text pairs, the proposed HCM module establishes the intricate associations within heterogeneous modalities of image-text pairs of the same person through HCM. It should be noted that PMM and HCM need to be combined to work properly. HCM comprises two components: PCM and ICPM. As shown in Table 4, No.0 vs No.2, No.0 vs No.3, and No.1 vs No.6, we readily observe that the inclusion of the PMM and HCM proves beneficial for enhancing the model’s performance. Removing this module will result in a certain degree of performance degradation.
Outlier Pseudo Label Mining (OPLM)
@@
As shown in Table 4, No.0 vs No.1, No.2 vs No.6 and No.3 vs No.5, we conducted an ablation study of OPLM. OPLM makes significant contributions to our method’s performance. Being a plug-and-play component, OPLM can seamlessly integrate into weakly supervised text-based person retrieval methods. These findings unequivocally demonstrate the effectiveness and flexibility of our OPLM approach. The removal of this module results in the inability to extract valuable diverse samples, leading to performance degradation of the model. We record the number of outliers during training and OPLM effectively reduces the number of outliers. Please refer to Appendix B for more detailed analysis.
cross-modal PCM vs single-modal PCM
We tried two types of prototypical contrast matching (PCM): cross-modal PCM and single-modal PCM, denoted as $PCM_{c}$ and $PCM_{s}$. In our all experiments, the default is $PCM_{c}$. The experimental results are shown No.0 vs No.3 in Table 5 and $PCM_{s}$ achieved a competitive performance on CUHK-PEDES. But $PCM_{c}$ surpassed $PCM_{s}$ in terms of all metrics and get better performance on CUHK-PEDES. $PCM_{c}$ incorporates both intra-modal and inter-modal prototype-based contrastive matching losses and enables more comprehensive representation learning across and within modalities, which is verified by visualization of t-SNEAppendix C.
Two-stage OPLM vs One-stage OPLM
Although we have developed a One-stage OPLM capable of mining valuable unclustered samples, certain image-text pairs remain unminable. Recognizing this limitation, we introduce a Two-stage OPLM to fully leverage these unminable image-text pairs. We conduct experiments to assess the effectiveness of this two-stage training scheme. As evidenced by the comparison between No.3 vs No.2 in Table 5, CPCL demonstrates superior performance when trained with the Two-stage OPLM scheme.
Exploration of Momentum updating scheme for PMM
PMM adopts a momentum updating scheme to dynamically refine the prototypical features. The performance gap between No.1 and No.3 in Table 5 demonstrates the importance of this scheme and its positive effect on CPCL. The dynamic momentum updating scheme offers the advantage of providing more confident learning targets. However, this approach also carries inherent risks, such as the potential for high variance and instability in the training process. Interestingly, during our experiments, we obviously did not observe these adverse phenomena. This suggests that the benefits of dynamic updating may outweigh its risks in certain contexts.
No. Methods CUHK-PEDES
R@1 R@5 R@10
0 CPCL+Kmeans-5000 68.68 86.26 91.16
1 CPCL+Kmeans-8000 69.80 86.36 91.11
2 CPCL+Kmeans-11000 70.16 86.84 91.89
3 CPCL+Kmeans-14000 69.86 86.24 91.20
5 CPCL(DBSCAN) 70.03 87.28 91.78
Table 6: Evaluate CPCL with other clustering algorithms on CUHK-PEDES.
Analysis of clustering algorithm
In order to verify that CPCL is still effective with other clustering algorithms, we conduct experiments by replacing the DBSCAN algorithm with $K$-means clustering algorithms where the $K$ is set as 5000, 8000, 11000 and 14000 for CUHK-PEDES. As shown in Table 6, we observe that the $K$ value impacts the model performance. The closer the $K$ value is to the actual identity number, the better the performance is as shown in No.0, No.1, and No.2 in Table 6. The performance of CPCL+Kmeans-11000 (No.2) in Table 6 is better than that of CPCL(DBSCAN), and we attribute this result to the fact that the $K=11000$ almost equals to the actual identity number of CUHK-PEDES(11003).
No. Initialization policy CUHK-PEDES
R@1 R@5 R@10
0 Random 69.14 85.92 91.03
1 Average 70.03 87.28 91.78
Table 7: Ablation study on PCM initialization policy.
Analysis of Initialization of PMM
In Table 7, we evaluate different initialization policies Of PCM. Specifically, we explored two initialization methods: “Random Initialization” and “Average Initialization.” In the “Random Initialization” strategy, we randomly select the feature of an instance belonging to a person to initialize his/her prototypical feature. This approach is straightforward and simple, but it is susceptible to the randomness of a single instance’s feature, especially when the dataset contains noisy labels. The randomly chosen feature may not accurately represent the person’s overall characteristics. In contrast, the “Average Initialization” strategy involves averaging the features of all samples belonging to the same person to initialize his/her prototypical feature. This method has the advantage of better reflecting a person’s overall characteristics, thereby enhancing the robustness and reliability of the initialized feature.
5 Conclusion
In this paper, we propose CPCL, which demonstrates competitive performance by enhancing alignment through intra-modal and inter-modal prototype-based contrastive matching losses. CPCL not only advances the field of weakly supervised text-based person retrieval but also offers a promising foundation for future research in pseudo-label refinement and domain generalization. We hope that our work can inspire and facilitate further innovations in weakly supervised learning and multi-modal data integration. Despite the encouraging results, a noticeable performance gap remains when compared to fully supervised text-based person retrieval. We believe that the quality of pseudo labels is a critical factor contributing to this gap. Future work will focus on improving pseudo-label quality and exploring more robust alignment strategies to further narrow this gap.
Appendix A Appendix
In this appendix, we present additional experimental results and detailed analyses that could not be included in the main paper due to space limitations. Appendix C presents the qualitative results, such as the visualization of embeddings of images and text, visualization of Saliency map, and visulization of retrieval results. Section 3.4 shows the number of outliers during trianing, which demonstrates the effectiveness of our Outlier Pseudo Label Mining (OPLM) module. Appendix D offers the performance of our method in the domain generalization problem. Finally, Appendix E delves into the ifluence of clustering hyper-parameters on our method. Appendix F provides the implementations details in our experiments. Appendix G provides the details of CUHK-PEDES, ICFG-PEDES, and RSTPReid datasets.
Figure 4: Illustration of the dynamically changing image outlier number on CUHK-PEDES dataset.
Figure 5: Illustration of the dynamically changing text outlier number on CUHK-PEDES dataset
Appendix B The Number of Outliers during Training
To maximize the utility of existing image-text pairing information, we introduce an Outlier Pseudo Label Mining (OPLM) module. This module is designed to further distinguish valuable outlier samples from each modality by mining implicit relationships between image-text pairs, thereby creating more reliable clusters.
We illustrate the quantity of image and text outliers during the initial 20 epochs of training on the CUHK-PEDES dataset in Fig. 4 and Fig. 5. The figures clearly show that the quantity of image and text outliers is significantly reduced when the OPLM module is employed, compared to the scenario without the OPLM module.
This reduction in outlier quantity demonstrates the efficacy of our OPLM in identifying and leveraging valuable outlier samples through the implicit relationships between image-text pairs. By effectively mining these relationships, the OPLM module enhances the clustering process, leading to more accurate and reliable clusters.
Figure 6: t-SNE visualization of random 10 person in CUHK-PEDES. Text and image samples are denoted as triangles and circles, respectively. Samples with the same color indicate that they belong to the same person. The same letter means that it corresponds to the same person.